OpenStudy (anonymous):

what are the roots of the quadratic equation x^2-2x-24=0

OpenStudy (anonymous):

-4 and 6

OpenStudy (anonymous):

Either use the quadratic formula, or factor. Let's try factoring: x^2 - 2x - 24 = 0 Now find a,b such that a + b = -2 and ab = -24 6 and 4 work here, one of them negative: (x - 6)(x + 4) = 0 Now for that to be true, what does x have to be?

OpenStudy (anonymous):

@rav is 4 negative and 6 positive

OpenStudy (anonymous):

I don't understand how to get x.

OpenStudy (anonymous):

We are setting (x - 6)(x + 4) = 0. That means that either (x - 6) = 0, or (x + 4) = 0 So for x - 6 = 0, we can use x = 6: x - 6 = 0 6 - 6 = 0 And we can use -4 for (x+4) = 0: x + 4 = 0 -4 + 4 = 0

OpenStudy (anonymous):

I understood better. thnk u